OpenStudy (anonymous):

use the identities: cos2x-cos^2x=0 2sinxtanx-tanx=1-2sinx 4sin^2x+2cos^2x=3 please help me! please

OpenStudy (tukitw):

Are you trying to solve for x?

OpenStudy (anonymous):

no i want the left side of the problem to equal whats on the right for the first one i need the left side to be solved so that the answer results in zero just like the right side

OpenStudy (tukitw):

It actually is what I meant

OpenStudy (anonymous):

oh right sorry i miss understood

OpenStudy (tukitw):

Did the question gives any range for x?

OpenStudy (anonymous):

no they dont

OpenStudy (tukitw):

Then I shall assume x has an infinite amount of solutions

OpenStudy (anonymous):

yeah

OpenStudy (tukitw):

cos (2x) - cos^2 x = 0 By the double-angle formula, cos (2x) = 2 cos^2 x - 1. 2 cos^2 x - 1 - cos^2 x = 0 cos^2 x - 1 = 0 cos^2 x = 1 cos x = 1 or cos x = -1 x = 2n pi or x = (2n + 1)pi where n is an element of the set of integers.

OpenStudy (tukitw):

I am starting to wonder if they require ambiguous values or specific values as answers

OpenStudy (anonymous):

the only thing it said was solve by using the identities

OpenStudy (tukitw):

Do you need help with factoring the expressions or solving for x or both?

OpenStudy (anonymous):

i do... so the first one it correct? with the second

OpenStudy (anonymous):

is there a way we can only solve by through identities and not factoing

OpenStudy (tukitw):

2 sin x tan x - tan x = 1 - 2 sin x tan x(2 sin x - 1) = 1 - 2 sin x tan x(2 sin x - 1) + (2 sin x - 1) = 0 (tan x + 1)(2 sin x - 1) = 0 tan x = -1 or sin x = 0.5

OpenStudy (tukitw):

Really, the second question seems easier to just factorise, in fact, I don't even know if using identities can help to solve this one

OpenStudy (anonymous):

could we change tan tp sin/cos

OpenStudy (tukitw):

Tried that and felt like I am straying off

OpenStudy (anonymous):

2sinx*sinx/cosx-sinx/cosx=....

OpenStudy (anonymous):

2sin^2x/cos-sinx/cosx

OpenStudy (anonymous):

sin^2x/cosx

OpenStudy (anonymous):

could it work out like this

OpenStudy (tukitw):

2 sin^2 x / cos x - sin x / cos x = 1 - 2 sin x 2 sin^2 x - sin x = cos x - 2 sin x cos x 2 sin^2 x - sin x = cos x - sin (2x) 1 - cos (2x) - sin x = cos x - sin (2x) It all seems weird to me

OpenStudy (anonymous):

your right i will stick with factoring... and the last one can you please help me with that one please

OpenStudy (tukitw):

4 sin^2 x + 2 cos^2 x = 3 4 sin^2 x + 2(1 - sin^2 x) = 3 4 sin^2 x + 2 - 2 sin^2 x = 3 2 sin^2 x = 1 sin^2 x = 0.5 sin x = sqrt 0.5 or sin x = -sqrt 0.5

OpenStudy (anonymous):

oh okay and you switched 2cos identity?

OpenStudy (tukitw):

You should probably just write the identities on a piece of paper

OpenStudy (tukitw):

sin^2 x + cos^2 x = 1 cos^2 x = 1 - sin^2 x

OpenStudy (anonymous):

im doing that right now...thank you so much!